OpenStudy (anonymous):

The formula for the area A of a trapezoid is A+ 1/2(b1-b2)h, where b1 and b2 represent the lengths of the bases and h represents the height. What equation can be used to find the height of a trapezoid?

Directrix (directrix):

A= 1/2*(b1-b2)*h Multiply both sides by 2 2A = 2 * (1/2) (b1 + b2)*h 2A = (b1 + b2)*h Divide by sides by (b1 + b2) 2A / (b1+b2) = h